# Are Ensembles Getting Better All the Time?

### v2

@@ 2.2.2 How to aggregate predictions?
Perhaps the simplest idea to aggregate predictions then to take their (unweighted) mean, namely to consider the prediction
@@
Several other aggregation schemes can, in some settings, be equivalent to the simple average. For instance, the geometric mean can be written as a regular mean in log scale, since $K]{\hat{y}_{1}\cdots\hat{y}_{K}}=\exp\left(\frac{1}{K}\sum_{k=1}^{K}\log\hat{y}_{k}\right)$. The popular method of majority voting can also be seen as a form of standard average. Indeed, consider for instance a binary classification problem for which several votes $\hat{y}_{1},\ldots,\hat{y}_{K}\in\{0,1\}$ are cast for a true class $y\in\{0,1\}$. The prediction space $\{0,1\}$ is nonconvex, but we may choose to see the votes as elements of its convex hull $C=[0,1]$. A natural choice of loss function is then the classification error $L(\hat{y})$, defined by Eq. (2). The most natural way of combining these votes is to use majority voting, considering as final choice $\tilde{y}_{\text{MV,$K$}}=\mathbf{1}(\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}>0.5)$. It turns out this majority vote will have the same loss as the ensemble prediction: $L(\tilde{y}_{\text{MV,$K$}})=L(\overline{y}_{K})$, which means that majority voting is equivalent to our simple average ensemble.
@@ 2.2.3 What sort of results are we going to prove?
A few comments about what this means are in order. The quantity $\mathbb{E}[L(\overline{y}_{K})]$ corresponds to the loss of a single prediction problem, e.g. predicting the label of a single image, averaged over the randomness of the ensemble. This is in contrast with many machine learning results that average over a dataset. Nevertheless, the monotonicity of individual losses will imply monotonicity for losses averaged over datasets by simply applying our results to every single point of the dataset and using the fact that an average of decreasing (respectively increasing) sequences remains decreasing (respectively increasing). This is illustrated in Fig. 2: our results imply that the cross-entropy loss decreases for every single test point, and in turn that the whole test cross-entropy decreases. Things are slightly different in Fig. 1: we cannot average our results over the whole test set (since it is composed of points whose accuracy eventually grows and points whose accuracy eventually decreases). However, we can split the test dataset into two sub-datasets for which we can apply our theorems, leading to the middle and right panels of Fig. 1.
@@ 3.1 Warm-up: An ensemble predicts better than a single model
Let us begin by reviewing a few classical arguments in favour of ensembling with a convex loss, based on Jensen’s inequality. Similar arguments were for instance given by McNees (1992) or Breiman (1996, Section 4) in the specific cases of the squared and absolute losses and by Perrone (1993) and George (1999) for a generic convex loss. The convexity of $L$ directly implies that, for all $\hat{y}_{1},\ldots,\hat{y}_{K}\in C$,
@@
When all elements of the ensemble are identically distributed, so are $L(\hat{y}_{1}),\ldots L(\hat{y}_{K})$ and integrating the left-hand side of Eq. (4) leads to
@@ 3.2 A first monotonicity lemma
$\begin{bmatrix}1\\ 1\\ \vdots\\ 1\\ 1\end{bmatrix}$ $=\frac{1}{K-1}\left(\begin{bmatrix}0\\ 1\\ \vdots\\ 1\\ 1\end{bmatrix}+\begin{bmatrix}1\\ 0\\ \vdots\\ 1\\ 1\end{bmatrix}+\cdots+\begin{bmatrix}1\\ 1\\ \vdots\\ 0\\ 1\end{bmatrix}+\begin{bmatrix}1\\ 1\\ \vdots\\ 1\\ 0\end{bmatrix}\right).$ (32)
@@
This inequality may already be interpreted as a monotonicity result: removing an element of the ensemble uniformly at random will hurt, on average. Note that this comes without any assumption on the ensemble itself, which could be composed of very different and potentially deterministic models. To prove stronger results, we need to make additional assumptions on the ensemble. Our assumptions will be probabilistic, and will translate the fact that we suspect a priori that all elements of the ensembles perform as well.
@@ 3.3 Exchangeable ensembles are getting better all the time
We saw in the previous section (Eq. (6)) that assuming that the ensemble is identically distributed led to an asymptotic form of monotonicity: the infinite ensemble $\overline{y}_{\infty}$ performs better than any finite ensemble, which in turn performs better than any single model, because of Eq. (5). However, the identically distributed assumption alone will not take us very far. Indeed, let us consider the following example: an ensemble of three models such that $\hat{y}_{1}$ and $\hat{y}_{2}$ are independent and identically distributed, and $\hat{y}_{3}=\hat{y}_{1}$. Then, one can show that the ensemble gets worse in average when adding its third component:
$\mathbb{E}\left[L\left(\frac{\hat{y}_{1}+\hat{y}_{2}}{2}\right)\right]\leq\mathbb{E}\left[L\left(\frac{\hat{y}_{1}+\hat{y}_{2}+\hat{y}_{3}}{3}\right)\right]$ (35)
@@
where $=_{d}$ denotes equality in distribution.
@@
However, there are also ensembles for which the exchangeability assumption is inherently wrong. This is notably the case of boosting, or any similar technique that trains models iteratively, making the construction of the $k$th base model dependent on the $k-1$ previous models, and hereby making their ordering essential. This is also the case when we have prior knowledge that some models are much better than others. In these two cases, using a weighted average instead of just the usual mean $\overline{y}_{K}$ is more advisable.
@@
$\small\mathbb{E}\left[L\left(\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}\right)\right]\leq\frac{1}{K}\sum_{j=1}^{K}\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k\neq j}\hat{y}_{k}\right)\right]\,.$ (40)
@@
$\small\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k\neq j}\hat{y}_{k}\right)\right]=\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k=1}^{K-1}\hat{y}_{k}\right)\right]\,,$ (41)
@@ 3.4 Strict improvements for strongly convex losses
[…]
@@ 4.1 Smooth nonconvex losses
A first case of nonconvex loss that is easy to tackle is the one of a concave loss. Indeed, if $L$ is concave, then $-L$ is convex which, by virtue of Theorem 2, implies that $\mathbb{E}[-L(\overline{y}_{K})]$ decreases and thus that $\mathbb{E}[L(\overline{y}_{K})]$ increases: for a concave loss, ensembles are getting worse all the time! However, to the best of our knowledge, concave losses are never used in statistics and machine learning. Laan et al. (2017) considered the example of a “square-root cost” $L(\hat{y})=\sqrt{|y-\hat{y}|}$ which is locally concave everywhere except at zero, but not globally concave (which prevents us from using Theorem 2).
@@ 4.1.1 Examples of smooth nonconvex losses
  • Scoring rules are the most common way to design losses in the probabilistic setting (when $\hat{y}\in[0,1]$). Most strictly proper scoring rules for classification are convex, with the notable exception of the spherical score (see, e.g., Gneiting and Raftery, 2007). Like the sigmoid loss, the spherical score is convex when $\hat{y}$ is small and concave when $\hat{y}$ is large. However, unlike the sigmoid whose inflection point was at $\hat{y}=50\%$, the spherical inflection point of is at $\hat{y}=1/8+\sqrt{17}/8\approx 64\%$. We give a formal definition in Appendix B. For the spherical loss, a “right prediction” is thus one that will predict the true label $y=0$ using the somewhat peculiar cutoff of $64\%$ instead of the usual $50\%$. This illustrates the asymmetry of the spherical score.
  • One way to avoid overfitting is to use a loss function that makes the model pay a price for overconfident predictions (potentially even correct ones). This is the rationale behind the tangent loss, introduced by Masnadi-Shirazi et al. (2010) as a boosting objective. Since this is not a probabilistic loss, predictions are real-valued scores $\hat{y}\in\mathbb{R}$. The tangent loss is concave when $\hat{y}$ is either too large or to small, penalising both overconfident mistakes and successes. In this case, “right predictions” correspond to the ones that are not overconfident.
Regarding regression, the most commonly occurring nonconvex smooth losses are used to induce robustness to outliers. The main motivation for such losses is that the squared loss will typically make the model pay an extremely large price for large errors. A way to fix this behavior is to design losses that resemble the squared error when errors are small, but grow less steadily when errors are large. The most famous robust loss is Huber’s (1964), which is equal to the squared error when the error is below a threshold, and grows linearly when the error is above this threshold. Huber’s loss is convex, therefore the analysis of the previous section guarantees that ensembles will get better all the time. However, many other losses that follow this rationale are nonconvex. A nice overview of these robust regression losses was recently offered by Barron (2019), who introduced a new general loss that generalises many previous losses (both convex and nonconvex). All nonconvex losses encompassed by Barron’s generalisation (as well as others, such as using a Student’s $t$ likelihood) share the property of being convex when the error $(\hat{y}-y)^{2}$ is smaller than a prespecified threshold $c$, and nonconvex when the error is larger than $c$. For all these nonconvex losses “bad predictions” are therefore outliers and “good” ones correspond to inliers. A few of these robust losses are displayed on Figure 3, all of them were chosen in order to have a cutoff of $c=10$.
@@ 4.1.2 Monotonicity of smooth nonconvex losses
Does this result says something about the classification error?
@@ 4.2.2 A deceivingly simple counter-example
Consider a binary classification classification task, where the true label is 00, and our predictions $\hat{y}_{1},\cdots,\hat{y}_{K}$ are i.i.d. samples from a Bernoulli distribution $\mathcal{B}(\overline{y}_{\infty})$. In that case, the classification error of any $\hat{y}\in[0,1]$ is simply $L(\hat{y})=\mathbf{1}(\hat{y}\leq 0.5)$. The expected loss of the ensemble will therefore be $\mathbb{E}[L(\overline{y}_{k})]=\mathbb{P}(\overline{y}_{K}\leq 0.5)$. It seems reasonable to conjecture that $\mathbb{P}(\overline{y}_{K}\leq 0.5)$ will go to zero decreasingly when $\overline{y}_{\infty}>0.5$, and to one increasingly when $\overline{y}_{\infty}<0.5$, which would be consistent with the analogy to the sigmoid loss sketched above. Unfortunately, this is not the case, and $\mathbb{P}(\overline{y}_{K}\leq 0.5)$ is not, in general, monotonic (even asymptotically).
Indeed, let us focus for instance on the case where the ensemble is eventually right, i.e., when $\overline{y}_{\infty}<0.5$. In that case, it can be shown that the subsequence of ensembles containing an odd number of models $L(\overline{y}_{2K+1})$ decreases (see Shteingart et al., 2020, Theorem 2, for a proof). However, monotonicity is broken when we include even numbers of models (Lam and Suen, 1997). A heuristic explanation of this failure, already mentioned by Condorcet (1785), is that odd numbers of predictions allow ties, while even numbers do not. If we alter slightly the problem to always allow ties by allowing jurors to abstain from predicting, then monotonicity eventually holds, as we discuss in the next section and in Appendix E. Probst and Boulesteix (2018, Section 3.1.1) discuss this problem of ties as well, and possible tie-breakers.
@@ 4.2.3 A useful formulation of the expected classification error
$L(\hat{y})=\left\{\begin{matrix}1\;\text{ if }\;\hat{y}^{(1)}\leq\text{max}_{c\neq 1}\left(\hat{y}^{(c)}\right)\\ 0\;\text{ if }\;\hat{y}^{(1)}>\text{max}_{c\neq 1}\left(\hat{y}^{(c)}\right).\end{matrix}\right.$ (47)
@@
[…]
$\mathbb{E}[L(\overline{y}_{K})]=1-\mathbb{P}\left(\overline{X}_{K}>0\right)=\mathbb{P}\left(-\overline{X}_{K}\geq 0\right).$ (48)
@@
Now that we have written the classification error as a simple tail probability, it seems natural to attack it using standard probabilistic asymptotics. As mentioned before, low-order asymptotics are not sufficient, and will will leverage “strong” large deviation theorems, that can be viewed as higher-order large deviations. This analysis is carried out in the next, and section is largely independent from the rest of the paper, since it may be of independent interest, beyond ensembles.

### v3

@@ 2.2.2 How to aggregate predictions?
Perhaps the simplest idea to aggregate predictions is to take their (unweighted) mean, namely to consider the prediction
@@
Several other aggregation schemes can, in some settings, be equivalent to the simple average. For instance, the geometric mean can be written as a regular mean in log scale, since $\sqrt[K]{\hat{y}_{1}\cdots\hat{y}_{K}}=\exp\left(\frac{1}{K}\sum_{k=1}^{K}\log\hat{y}_{k}\right)$. The popular method of majority voting can also be seen as a form of standard average. Indeed, consider for instance a binary classification problem for which several votes $\hat{y}_{1},\ldots,\hat{y}_{K}\in\{0,1\}$ are cast for a true class $y\in\{0,1\}$. The prediction space $\{0,1\}$ is nonconvex, but we may choose to see the votes as elements of its convex hull $C=[0,1]$. A natural choice of loss function is then the classification error L(y^)L($\hat${y}), defined by Eq. (2). The most natural way of combining these votes is to use majority voting, considering as final choice $\tilde{y}_{\text{MV,$K$}}=\mathbf{1}(\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}>0.5)$. It turns out this majority vote will have the same loss as the ensemble prediction: $L(\tilde{y}_{\text{MV,$K$}})=L(\overline{y}_{K})$, which means that majority voting is equivalent to simple averaging.
@@ 2.2.3 What sort of results are we going to prove?
A few comments about what this means are in order. The quantity $\mathbb{E}[L(\overline{y}_{K})]$ corresponds to the loss of a single prediction problem, e.g., predicting the label of a single image, averaged over the randomness of the ensemble. This is in contrast with many machine learning results that average over a dataset. Nevertheless, the monotonicity of individual losses will imply monotonicity for losses averaged over datasets by simply applying our results to every single point of the dataset and using the fact that an average of decreasing (respectively increasing) sequences remains decreasing (respectively increasing). This is illustrated in Fig. 2: our results imply that the cross-entropy loss decreases for every single test point, and in turn that the whole test cross-entropy decreases. Things are slightly different in Fig. 1: we cannot average our results over the whole test set (since it is composed of points whose accuracy eventually grows and points whose accuracy eventually decreases). However, we can split the test dataset into two sub-datasets for which we can apply our theorems, leading to the middle and right panels of Fig. 1.
@@ 3.1 Warm-up: An ensemble predicts better than a single model
Let us begin by reviewing a few classical arguments in favour of ensembling with a convex loss, based on Jensen’s inequality. The simple and compelling reasoning below, that dates back to early works on forecasting at least (Winkler, 1971, Section 3.4), was called the “algebraic wisdom of crowds” by Manski (2016). Similar arguments were also given, for instance, by McNees (1992) or Breiman (1996, Section 4) in the specific cases of the squared and absolute losses and by Perrone (1993) and George (1999) for a generic convex loss. The convexity of $L$ directly implies that, for all $\hat{y}_{1},\ldots,\hat{y}_{K}\in C$,
@@
When all elements of the ensemble are identically distributed, so are $L(\hat{y}_{1}),\ldots,L(\hat{y}_{K})$ and integrating the left-hand side of Eq. (4) leads to
@@ 3.2 A first monotonicity lemma
[…]
@@
This inequality may already be interpreted as a monotonicity result: removing an element of the ensemble uniformly at random will hurt, on average. Note that this comes without any assumption on the ensemble itself, which could be composed of very different and potentially deterministic models. To prove stronger results, we need to make additional assumptions on the ensemble. Our assumptions will be probabilistic, either requiring independent, identically-distributed predictions, or a weakened version of this assumption called exchangeability. They will translate the fact that we suspect that, a priori, all elements of the ensembles perform equally well.
@@ 3.3 Exchangeable ensembles are getting better all the time
We saw in the previous section (Eq. 6) that assuming that the ensemble is identically distributed led to an asymptotic form of monotonicity: the infinite ensemble $\overline{y}_{\infty}$ performs better than any finite ensemble, which in turn performs better than any single model, because of Eq. (5). However, the identically distributed assumption alone will not take us very far. Indeed, let us consider the following example: an ensemble of three models such that $\hat{y}_{1}$ and $\hat{y}_{2}$ are independent and identically distributed, and $\hat{y}_{3}=\hat{y}_{1}$. Then, one can show that the ensemble gets worse in average when adding its third component:
$\mathbb{E}\left[L\left(\frac{\hat{y}_{1}+\hat{y}_{2}}{2}\right)\right]\leq\mathbb{E}\left[L\left(\frac{\hat{y}_{1}+\hat{y}_{2}+\hat{y}_{3}}{3}\right)\right]\,.$ (10)
@@
where $_{d}$ denotes equality in distribution.
@@
However, there are also ensembles for which the exchangeability assumption is inherently wrong. This is notably the case of boosting, or any similar technique that trains models iteratively, making the construction of the $k$th base model dependent on the k1k-1 previous models, and hereby making their ordering essential. This is also the case when we have prior knowledge that some models are much better than others. In these two cases, using a weighted average instead of just the usual mean $\overline{y}_{K}$ is more advisable.
@@
$\mathbb{E}\left[L\left(\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}\right)\right]\leq\frac{1}{K}\sum_{j=1}^{K}\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k\neq j}\hat{y}_{k}\right)\right]\,.$ (15)
@@
$\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k\neq j}\hat{y}_{k}\right)\right]=\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k=1}^{K-1}\hat{y}_{k}\right)\right]\,$, (16)
@@ 3.4 Strict improvements for strongly convex losses
The previous result show that the average loss is non-increasing. Under which conditions is it strictly decreasing? This depends both on the properties of the loss (for instance, if $L$ is constant, no ensemble can strictly improve) and of the ensemble (an ensemble containing identical copies of the same predictor will have constant loss). We will see now that two simple assumptions are sufficient to ensure strict improvements: strong convexity of the loss, and existence of a nonsingular covariance of the ensemble. In this subsection, we further assume that the prediction set $C$ is equipped with a norm $\|\cdot\|$.
@@ 4.1 Smooth nonconvex losses
A first case of nonconvex loss that is easy to tackle is the one of a concave loss. Indeed, if $L$ is concave, then L-L is convex which, by virtue of Theorem 2, implies that $\mathbb{E}[-L(\overline{y}_{K})]$ decreases and thus that $\mathbb{E}[L(\overline{y}_{K})]$ increases: for a concave loss, ensembles are getting worse all the time! However, to the best of our knowledge, concave losses are never used in statistics and machine learning. Laan et al. (2017) considered the example of a “square-root cost” $L(\hat{y})=\sqrt{|y-\hat{y}|}$ which is locally concave everywhere except at zero, but not globally concave (which prevents us from using Theorem 2).
@@ 4.1.1 Examples of smooth nonconvex losses
  • Scoring rules are the most common way to design losses in the probabilistic setting (when $\hat{y}\in[0,1]$). Most strictly proper scoring rules for classification are convex (e.g. the cross-entropy or the Brier score), with the notable exception of the spherical score (see, e.g., Gneiting and Raftery, 2007). Like the sigmoid loss, the spherical score is convex when $\hat{y}$ is small and concave when $\hat{y}$ is large. However, unlike the sigmoid whose inflection point was at $\hat{y}=50\%$, the inflection point of the spherical score is at $\hat{y}=1/8+\sqrt{17}/8\approx 64\%$. We give a formal definition in Appendix B. For the spherical loss, a “right prediction” is thus one that will predict the true label $y=0$ using the somewhat peculiar cutoff of $64\%$ instead of the usual $50\%$. This illustrates the asymmetry of the spherical score.
  • One way to avoid overfitting is to use a loss function that makes the model pay a price for overconfident predictions (potentially even correct ones). This is the rationale behind the tangent loss, introduced by Masnadi-Shirazi et al. (2010) as a boosting objective. Since this is not a probabilistic loss, predictions are real-valued scores $\hat{y}\in\mathbb{R}$. The tangent loss is concave when $\hat{y}$ is either too large or to small, penalising both overconfident mistakes and successes. In this case, “right predictions” correspond to the ones that are not overconfident. For more details on the tangent loss, see Appendix B.
Regarding regression, the most commonly occurring nonconvex smooth losses are used to induce robustness to outliers. The main motivation for such losses is that the squared loss will typically make the model pay an extremely large price for large errors. A way to fix this behavior is to design losses that resemble the squared error when errors are small, but grow less steadily when errors are large. The most famous robust loss is Huber’s (1964), which is equal to the squared error when the error is below a threshold, and grows linearly when the error is above this threshold. Huber’s loss is convex, therefore the analysis of the previous section guarantees that ensembles will get better all the time. However, many other losses that follow this rationale are nonconvex. A nice overview of these robust regression losses was recently offered by Barron (2019), who introduced a new general loss that generalises many previous losses (both convex and nonconvex). All nonconvex losses encompassed by Barron’s generalisation (as well as others, such as using a Student’s $t$ likelihood) share the property of being convex when the error $(\hat{y}-y)^{2}$ is smaller than a prespecified threshold $c$, and nonconvex when the error is larger than $c$. For all these nonconvex losses “bad predictions” are therefore outliers and “good” ones correspond to inliers. A few of these robust losses are displayed on the right-hand side of Figure 3, all of them were chosen in order to have a cutoff of $c=10$.
@@ 4.1.2 Monotonicity of smooth nonconvex losses
Does this result say something about the classification error?
@@ 4.2.2 A deceivingly simple counter-example
Consider a binary classification task, where the true label is 0, and our predictions $\hat{y}_{1},\cdots,\hat{y}_{K}$ are i.i.d. samples from a Bernoulli distribution $\mathcal{B}(\overline{y}_{\infty})$. In that case, the classification error of any $\hat{y}\in[0,1]$ is simply $L(\hat{y})=\mathbf{1}(\hat{y}\geq 0.5)$. The expected loss of the ensemble will therefore be $\mathbb{E}[L(\overline{y}_{k})]=\mathbb{P}(\overline{y}_{K}\geq 0.5)$. It seems reasonable to conjecture that $\mathbb{P}(\overline{y}_{K}\geq 0.5)$ will go to zero decreasingly when $\overline{y}_{\infty}<0.5$, and to one increasingly when $\overline{y}_{\infty}>0.5$, which would be consistent with the analogy to the sigmoid loss sketched above. Unfortunately, this is not the case, and $\mathbb{P}(\overline{y}_{K}\geq 0.5)$ is not, in general, monotonic (even asymptotically).
Indeed, let us focus for instance on the case where the ensemble is eventually right, i.e., when $\overline{y}_{\infty}<0.5$. In that case, it can be shown that the subsequence of ensembles containing an odd number of models $L(\overline{y}_{2K+1})$ decreases (see Shteingart et al., 2020, Theorem 2, for a proof). However, monotonicity is broken when we include even numbers of models (Lam and Suen, 1997). A heuristic explanation of this failure, already mentioned by Condorcet (1785), is that even numbers of predictions allow ties, while odd numbers do not. If we alter slightly the problem to always allow ties by allowing jurors to abstain from predicting, then monotonicity eventually holds, as we discuss in the next section and in Appendix E. Probst and Boulesteix (2018, Section 3.1.1) discuss this problem of ties as well, and possible tie-breakers.
@@ 4.2.3 A useful formulation of the expected classification error
[…]
@@
For all $k\in\{1,\ldots,K\}$, let $X_{k}=\left({\hat{y}_{k}}^{(1)}-{\hat{y}_{k}}^{(2)},\ldots,{\hat{y}_{k}}^{(1)}-{\hat{y}_{k}}^{(n_{\text{Cl}})}\right)\in\mathbb{R}^{n_{\text{Cl}}-1}$ and $\overline{X}_{K}=\left({\overline{y}_{k}}^{(1)}-{\overline{y}_{k}}^{(2)},\ldots,{\overline{y}_{k}}^{(1)}-{\overline{y}_{k}}^{(n_{\text{Cl}})}\right)\in\mathbb{R}^{n_{\text{Cl}}-1}$. Then,
where the inequality inside the probability is to be taken entry-wise.
@@
Now that we have written the classification error as a simple tail probability, it seems natural to attack it using standard probabilistic asymptotics. As mentioned before, low-order asymptotics are not sufficient, and we will leverage “strong” large deviation theorems, that can be viewed as higher-order large deviations. This analysis is carried out in the next section, which is largely independent from the rest of the paper. We note that it may be of independent interest, beyond ensembles.